A 200 kg cart is pushed slowly up an incline how much work does the pushing force do in moving the cart up to a platform. 1.5 m above the starting point if the distance along the incline to the platform is 7 m and a friction force of 150 N

1
Expert's answer
2021-11-07T19:27:19-0500

"W=mgh+F\cdot l=200\cdot9.8\cdot 1.5+150\cdot7=3990 \ (N)"